Solve the system of linear equations by graphing
2x-y=0
4x+y=6

Answer by jim_thompson5910(35256)

Add the two equations to get

2x-y=0
4x+y=6
------------
6x+0y = 6

6x = 6

x = 6/6

x = 1

Now use this solution of x to find the solution for y. Let me know if this makes sense or not. Also, let me know if you still need help. Thanks.


--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------